# Exploring the Zero-Shot Capabilities of LLMs Handling Multiple Problems at once

### v2

@@ Abstract
Recent studies have proposed placing multiple problems in a single prompt to improve input token utilization for a more efficient LLM inference. We call this multi-problem prompting (MPP), in contrast to conventional single-problem prompting (SPP) that prompts an LLM with a single problem at a time. While MPP has been shown to work comparably well or even better than SPP under few-shot settings, its zero-shot performance is underexplored, which better reveals the innate multiple problem handling capabilities of LLMs. To address that, we study the zero-shot MPP performance of various LLMs on 6 classification and 12 reasoning benchmarks and confirm that LLMs are competent zero-shot multi-problem solvers. We also examine the conditions of effectiveness of zero-shot MPP and explore several model-level factors that may enable MPP. We observe that LLMs consistently perform worse with selecting indices of texts of a given class label and with multiple mixed-source reasoning problems. This indicates a lack of true understanding. We also find that instruction tuning is an important factor than enhances MPP.
@@
Exploring the Zero-Shot Capabilities of LLMs
Handling Multiple Problems at once
@@ 1 Introduction
Large language models (LLMs) have recently shown remarkable natural language understanding and reasoning capabilities (OpenAI, 2023; Gemini-Team, 2023; Anthropic, 2024), both under zero-shot (Wei et al., 2022; Kojima et al., 2022) and few-shot (Brown et al., 2020; Wei et al., 2023) settings. Given the non-trivial cost of deploying LLMs at large scale, substantial efforts have been made to make LLM inferences more cost efficient, e.g., distilling a smaller but highly capable model from a large one Abdin et al. (2024); OpenAI (2024).
@@
Besides model-level distillation, recent studies Cheng et al. (2023); Lin et al. (2024); Son et al. (2024) have proposed various prompt-level approaches that place multiple problems in a single prompt, which we collectively call multi-problem prompting (MPP). As illustrated in Fig 1, MPP (e.g., BatchClf) inserts multiple problems $P=(P_{1},P_{2},...,P_{n})$ after a shared context $C$ (e.g., task instruction and/or exemplars) to avoid repeating $C$ $n$ times for each problem as in standard single-problem prompting (SPP, e.g., SingleClf), which
Figure 1: Accuracy of 7 LLMs on 4 task types constructed from SST-2 with up to 5 problems, including standard zero-shot prompting (SingleClf), zero-shot multi-problem prompting (BatchClf), and its reformulations (SelectOne and SelectAll). The text boxes provide simplified task illustrations. See Section 3 for details.
improves input token utilization. In practice, MPP is proposed in few-shot settings, where the context, which contains exemplars, is typically significantly longer than individual problems and an increased inference efficiency is guaranteed for the less repetitions of the shared context Cheng et al. (2023). Empirically, MPP is also found to perform comparably well or even better than SPP across a wide range of domains under few-shot settings Cheng et al. (2023); Lin et al. (2024); Son et al. (2024).
Despite progress in few-shot MPP, the performance of LLMs in zero-shot MPP remains underexplored. This is significant on at least two levels. First, on a practical level, many real-world tasks, such as classification, are typically approached in zero-shot settings Ziems et al. (2024). Moreover, designing few-shot exemplars can be tedious and costly to obtain in practice Kojima et al. (2022); Yasunaga et al. (2024). Second, from a scientific perspective, zero-shot MPP provides deeper insights into the innate capabilities of LLMs concurrently handling multiple tasks, a key indicator of intelligence Meyer and Kieras (1997).
To this end, we evaluate LLMs from 4 model families with zero-shot MPP on multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks. To further understand the conditions of effectiveness of MPP, we also perform MPP with two task reformulations (i.e., SelectOne/SelectAll in Fig 1) and with multiple mixed-source problems leveraging these classification and reasoning benchmarks, respectively. While MPP is meant to be an inference-efficient prompting technique, the main motivation of the paper is to provide scientific understandings of LLM capabilities related to the concurrent handling of multiple problems through comprehensive and systematic contrastive evaluation. For this reason, we also explore additional evidence of strong multiple problem handling capabilities of LLMs and possible model-level factors that may enable MPP.
@@
  • We demonstrate the general effectiveness of zero-shot MPP, when the problems are homogeneous or from a single source.
  • We identify two general conditions under which all tested LLMs perform significantly worse than expected and explore the roles of several model-level factors in MPP.
  • We design a new benchmark ZeMPEB (Zero-shot Multi-Problem Evaluation Benchmark), comprising 53,100 zero-shot multi-problem prompts containing classification and reasoning problems, for evaluating the zero-shot multiple problem handling capabilities of LLMs. ZeMPEB will be released upon acceptance.
The paper proceeds as follows. We reviews related works in Section 2. We perform zero-shot MPP and related experiments with classification and reasoning problems in Sections 3 and 4, respectively. We separate the classification and reasoning experiments in two sections due to the different natures of classification and reasoning, which we leverage for more nuanced evaluations. We explore several model-level factors that may enable MPP in Section 5. We conclude in Section 6.
@@ 2 Related Work
Recently, Cheng et al. (2023) propose few-shot MPP named batch prompting that prompts LLMs with problems batched together from single sources following a few batches of equally sized exemplars. They found that few-shot MPP greatly increases LLM inference efficiency while retaining downstream performance with a small batch size (e.g., <6). To ensure that batch prompting works with large batch sizes, Lin et al. (2024) introduce a sampling optimization method that takes a majority vote over repeated permutations of batch samples.
@@
The related studies surveyed so far imply that LLMs have strong multiple problem handling capabilities under few-shot settings. To the best of our knowledge, Laskar et al. (2023) is the only study that shows instruction-tuned GPTs can handle 5 short questions sampled from two open-domain QA benchmarks at once zero-shot, but they do not conduct no further experiments. Overall, it is poorly understood how well LLMs can concurrently handle multiple problems zero-shot.
3 MPP: Classification
We first start our MPP experiments with classification problems, followed by experiments with reasoning problems in the next section.
@@ 3 MPP: Classification
Table 1: Benchmarks used in Section 3 (Classification) and Section 4 (Reasoning). We use the test splits wherever possible, except for CoLA, StrategyQA, and CommonsenseQA, for which we use the dev splits, since the test splits are not publicly available. For AGNews and SNLI, we randomly sample 1,000 examples from the test splits.
3.1 Data
@@
We consider the following 6 classification benchmarks, as described and referenced in Table 1: SST-2, CoLA, AGNews, MRPC, SNLI, and WiC. These benchmarks are commonly used for NLP evaluation, with SST-2, CoLA, and MRPC appearing in GLUE Wang et al. (2019) and WiC in SuperGLUE (Sarlin et al., 2020). They cover (1) two classification paradigms, i.e., single-text and text-pair classification, and (2) six distinct task objectives.
3.2 Evaluation Tasks
@@
For each benchmark with $m$ class labels $L=(L_{1},...,L_{m})$, we perform standard zero-shot single-problem prompting (SPP) on all the problems to establish baselines for LLMs. We randomly sample 100 distinct sets of 100 problems $\{{P^{(i)}}=(P_{1}^{(i)},P_{2}^{(i)},...,P_{100}^{(i)})\}_{i=1}^{100}$ out of each benchmark and take the first $n$ problems from each problem set $P^{(i)}$ to generate a multi-problem prompt of task size $n$, with the task size for SPP being 1. We consider the following 5 task sizes and ensure that each task size has 100 distinct problem sets: 5/10/20/50/100 for single-text benchmarks and 3/5/10/20/50 for text-pair benchmarks. Note that by design a larger task size only differs from a smaller one by having additional problems given a benchmark, which helps us isolate the effect of task size. In total, this results in 13,500 prompts for all MPP tasks.
As illustrated in Fig 1, there are a total of 4 classification-related task types. These task types are designed to minimally differ from one another such that model performance on a task pair with minimal differences immediately reveals its (lack of) ability in certain aspects we intend to test. We provide further descriptions of the 4 task types below, where every two neighboring task types make a minimally different task pair.
Single Classification (SingleClf)
This is SPP for classification. The prompt indicates the task objective as well as the possible class labels, followed by one target classification problem to be classified.
Batch Classification (BatchClf)
This is MPP for classification. The prompt for BatchClf is nearly identical to that of SingleClf to maximize comparability, expect that it incorporates multiple target classification problems to be classified at once.
Index Selection One Label (SelectOne)
SelectOne is a reformulation of BatchClf. Instead of doing batch classification, SelectOne instructs LLMs to select all indices out of $n$ problems $P$ belonging to each class label $L_{i}$ one at a time in $m$ separate prompts, i.e., $\{j:label(P_{j})=L_{i},j\in[1,n]\}$ for $i\in[1,m]$. The expected output for each $L_{i}$ is “None,” “All,” or a list of applicable indices.
Index Selection All Labels (SelectAll)
SelectAll prompts a system to do SelectOne for all possible class labels $L$ all at once. The expected output is thus a concatenation of SelectOne output for $L$, i.e., $\{(L_{i},\mathrm{SelectOne}(P,L_{i})):i\in[1,m]\}$.
3.3 Experimental Setups
LLMs and evaluation settings
@@ 3.3 Experimental Setups
We evaluate 7 LLMs from 4 model families with greedy decoding: Vicuna (13B, Chiang et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), Llama-3 8B and 70B (Instruct, Meta, 2024), GPT-3.5, and GPT-4 (OpenAI, 2023).
Performance metric
@@
We measure the average per-problem accuracy (PPA) to unify the evaluation across the four task types. PPA, defined in Equation 1, is the average accuracy of classifying $n$ problems in each prompt or, in the case of SelectOne, in each set of directly related prompts targeting different class labels.
@@
$I(P_{i})$ is the inferred LLM-generated answer to the $i$th problem in the input prompt, $A_{i}$ is the ground truth, and $\delta(i,j)=1$ iff $i=j$ and 0 otherwise. For the two index selection tasks, $I(P_{i})$ is determined by considering the LLM’s assignments of indices to all class labels. Other than assigning an index to a wrong class label, there are two more error types. First, LLMs may assign an index with more than one class label, i.e., an contradiction error. Second, LLMs may assign no labels to an index at all, namely, an non-excluded middle error.
@@
Figure 2: Aggregated average accuracy of the 7 LLMs on the 4 types of tasks across various task sizes for each benchmark. See Appendix B for the full results for each LLM.

### v3

# Evaluating LLMs with Multiple Problems at once

@@ Abstract
This paper shows the benefits and fruitfulness of evaluating LLMs with multiple problems at once, a paradigm we call multi-problem evaluation (MPE). Unlike conventional single-problem evaluation, where a prompt presents a single problem and expects one specific answer, MPE places multiple problems together in a single prompt and assesses how well an LLM answers all these problems in a single output. Leveraging 6 classification and 12 reasoning benchmarks that already exist, we introduce a new benchmark called ZeMPE (Zero-shot Multi-Problem Evaluation), comprising 53,100 zero-shot multi-problem prompts. We experiment with a total of 13 LLMs from 5 model families on ZeMPE to present a comprehensive and systematic MPE. Our results show that LLMs are capable of handling multiple problems from a single data source as well as handling them separately, but there are conditions this multiple problem handling capability falls short. In addition, we perform in-depth further analyses and explore model-level factors that may enable multiple problem handling capabilities in LLMs. We release our corpus and code [^1] https://github.com/jaaack-wang/multi-problem-eval-llm to facilitate future research.
@@
Evaluating LLMs with Multiple Problems at once
@@ 1 Introduction
Thanks to the advances in both GPU hardware and algorithms (Dai et al., 2019; Beltagy et al., 2020; Dao et al., 2022; Ding et al., 2024; Chen et al., 2024, inter alia), large language models (LLMs) have been developed with increasingly larger context windows (e.g., 8K, 128K, 2M). To leverage the extended context windows, recent studies Cheng et al. (2023); Lin et al. (2024); Son et al. (2024) have proposed various prompting strategies that place multiple problems in a single prompt, which we collectively call multi-problem prompting (MPP). The basic idea of MPP is to place multiple problems after a shared context $C$ (e.g., task instruction and/or exemplars) to avoid repeating $C$ for each problem as in standard single-problem prompting (SPP), which improves input token utilization and reduces LLM inference costs per problem.
In this study, we evaluate a wide range of LLMs with multiple problems at once through MPP, a paradigm we call multi-problem evaluation (MPE) Wang et al. (2025). [^2] While MPE is achieved through MPP, MPP can be used for purposes other than evaluation, e.g., knowledge retrieval, question answering, and other use cases. It it thus necessary to distinguish MPP from MPE and SPP from SPE. While the main goal of MPP is to improve the cost-efficiency of LLM inference, we view MPE primarily as a valuable evaluation paradigm for probing LLM capabilities, rather than merely a cost-saving engineering strategy. Unlike conventional single-problem evaluation that assesses an LLM’s ability to answer a single problem through SPP, MPE assesses an LLM’s ability to concurrently handle multiple problems at once or in a single output. Understanding the multiple problem handling capabilities of LLMs is an important research question because it gives us a foundational insight into how LLMs operate over multi-problem inputs that can be sufficiently long and use information from individual problems contained within each multi-problem input.
To enable a comprehensive and systematic MPE, we introduce ZeMPE (Zero-shot Multi-Problem Evaluation), a new benchmark comprising 53,1000 zero-shot multi-problem prompts. ZeMPE is synthetically generated by leveraging 6 classification and 12 reasoning benchmarks that already exist and are widely used. Moreover, ZeMPE includes various types of evaluation tasks to allow for deep and nuanced analyses, taking into account how multiple problems are presented in the prompt and whether these problems are sampled from the same data source or not. We do not mix classification and reasoning problems together due to the different natures of these two types of problems and to not make our experiments confounding.
@@
  • We show that LLMs are capable of handling multiple classification or reasoning problems from a single data source as well as handling them separately zero-shot. We present multiple pieces of evidence, in addition to direct performance comparisons to validate this.
  • We demonstrate that just like few-shot MPP, zero-shot MPP can be highly cost-efficient.
  • We identify two general conditions under which LLMs perform significantly worse than expected when presented with multiple problems and explore the roles of several model-level factors that may enable their multiple problem handling capabilities.
  • We release a new MPE benchmark called ZeMPE to facilitate future MPE studies.
@@ 2 Related Work
Recently, Cheng et al. (2023) propose few-shot MPP named batch prompting that prompts LLMs with problems batched together from single sources following a few batches of equally sized exemplars. They find that few-shot MPP greatly increases LLM inference efficiency while retaining downstream performance with a small batch size (e.g., <6). To ensure that batch prompting works with large batch sizes, Lin et al. (2024) introduce a sampling optimization method that takes a majority vote over repeated permutations of batch samples.
@@
In addition to these few-shot studies, Laskar et al. (2023) shows that instruction-tuned GPTs can handle 5 short questions sampled from two open-domain QA benchmarks at once zero-shot, but the base GPT models can barely perform the task. To the best of our knowledge, Wang et al. (2025) present the first systematic evaluation of LLMs’ zero-shot ability to tackle multiple homogeneous classification problems drawn from six standard benchmarks. They show that, while LLMs can usually solve several such classifications in a single prompt with accuracy comparable to handling them one-by-one, their performance deteriorates sharply when the prompt instead asks them to return the indices of texts belonging to each class–a shortfall that remains consistent across models, prompting conditions, and experimental settings.
Building on top of Wang et al. (2025), this study examines a total of 13 LLMs on 18 existing benchmarks, including 12 reasoning benchmarks that are not part of Wang et al. (2025)’s evaluation. Besides from reaffirming Wang et al. (2025)’s finding that LLMs are capable of handling multiple problems from a single data source as well as handling them separately, we perform in-depth further analyses to both validate such capabilities and expose their limitations. Moreover, we explore model-level factors that may enable LLM’s strong multiple problem handling capabilities.
3 Multi-Problem Evaluation
This section compares single-problem evaluation (SPE) and multi-problem evaluation (MPE) and introduces ZeMPE, a new MPE benchmark.
3.1 SPE vs. MPE
SPE assesses an LLM’s ability to solve a type of problem by prompting the LLM with such a problem one at a time. In contrast, MPE places multiple problems together that can be of a same or different types and evaluates how well an LLM handles them. A simple example of a multi-problem task would bundle multiple classification or QA problems together and ask LLMs to solve them sequentially.
3.2 Benefits of MPE
MPE has at least three advantages over SPE.
Lesser Data Contamination Concerns
First, it is less likely for LLMs to encounter exact multi-problem prompts during pre-training because of the combinatory nature of constructing prompts from multiple problems. This helps mitigate a growing data contamination concern in modern large-scale pre-training (Jacovi et al., 2023; Sainz et al., 2023).
Improved Controllability and Interpretability of Evaluation
Second, since we can manipulate what kind of problems and how many problems to include, we know exactly which problem an LLM gets wrong or right across positions in the prompts. This enables us to construct a well controlled and easily interpretable evaluation.
High Feasibility and Adaptability
Third, our study demonstrates that leveraging the rich existing benchmarks to create a new multi-problem task is cheap, easy to implement, and highly adaptable. The most laborious component is the prompt design, which, once done, can easily be applied to a set of benchmarks with minimal adaptation.
@@ 3.2 Benefits of MPE
Table 1: Existing benchmarks we use to construct ZeMPE. We use the test splits wherever possible, except for CoLA, StrategyQA, and CommonsenseQA, for which we use the dev splits, since the test splits are not publicly available. For AGNews and SNLI, we randomly sample 1,000 examples from the test splits.
3.3 ZeMPE
We describe how we construct ZeMPE as well as how we evaluate LLMs on it.
3.3.1 Data
We use 6 classification and 12 reasoning benchmarks, as described and referenced in Table 1, to ensure a comprehensive and systematic evaluation.
@@ 3.3.1 Data
The classification benchmarks are commonly used for NLP evaluation, with SST-2, CoLA, and MRPC appearing in GLUE Wang et al. (2019) and WiC in SuperGLUE (Sarlin et al., 2020). They cover two classification paradigms (single-text and text-pair) and six distinct task objectives.
The 12 reasoning benchmarks are widely utilized in LLM evaluation (Kojima et al., 2022; Wei et al., 2023; Zhang et al., 2023, inter alia). These benchmarks test symbolic reasoning (Coin Flips & Last Letters), commonsense reasoning (StrategyQA, CommonsenseQA, Object tracking, & Bigbench date), and arithmetic reasoning (AQuA, SVAMP, GSM8K, MultiArith, AddSub, & SingleEq), and require three answer formats (Yes/No, multiple choice, and free-response).
3.3.2 Evaluation Tasks and Prompt Design
We separate the classification and reasoning problems, due to their different natures and to avoid confounding experiments, when designing multi-problem evaluation tasks.
Unlike previous related studies Cheng et al. (2023); Lin et al. (2024); Son et al. (2024) that evaluate LLMs on multi-problem prompts under few-shot settings, our evaluation tasks are all zero-shot, which are rather underexplored, as shown in Section 2. Moreover, zero-shot MPE is significant on at least two levels. First, on a practical level, many real-world tasks, such as classification, are typically approached in zero-shot settings Ziems et al. (2024). Moreover, designing few-shot exemplars can be tedious and costly to obtain in practice Kojima et al. (2022); Yasunaga et al. (2024). Second, from a scientific perspective, zero-shot MPP may provide deeper insights into the innate capabilities of LLMs concurrently handling multiple tasks.
The evaluation tasks are as follows with the full prompt templates for each task in Appendix E.
Classification-Related Tasks
We call the standard classification task via SPP Single Classification or SingleClf, which serves as a baseline to be compared with MPE tasks. When an LLM is prompted to solve multiple homogeneous classification problems through MPP, this task is known as Batch Classification or BatchClf. Index Selection One Label (SelectOne) and Index Selection All Labels (SelectAll) are two reformulations of BatchClf. Instead of making multiple classifications under BatchClf, these two tasks instruct LLMs to select indices of text falling into each class label, either independently in m separate prompts (SelectOne) or altogether in a single prompt (SelectAll), where m is the number of class labels in a benchmark.
We design the two selection tasks to test LLM’s understanding of the classifications performed under BatchClf. Since selection tasks of size $n$ may have anywhere from 0 to $n$ correct indices per class, spurious correlations are less likely during our evaluation, given the combinatory answer space.
For each of the four tasks above, we start by describing the task in the prompt and then include one or multiple classification problems afterwards. LLMs are instructed to solve these problems according to the specific task requirements.
Reasoning-Related Tasks
We first test on all the reasoning problems in each benchmark to establish LLM SPP baselines. Two MPE tasks are designed, i.e., single-source and mixed-source multi-problem reasoning, or MultiReasonSS and MultiReasonMS. For both tasks, we bundle multiple reasoning problems together with an indexed prefix “Qi: ,” where i is the index of each problem starting from “Q1.” We use two headers, namely, “Questions,” before the bundled questions, and “Answers,” before answers to be generated by LLMs. Unlike prompts designed for classification-related tasks, there is no shared task description, since each question already describes its own unique problem to be solved.
3.3.3 ZeMPE Composition
We define task size $n$ as the number of classification or reasoning problems included in a prompt. We construct a multi-problem prompt with all problems sampled from the same benchmark, except for MultiReasonMS where we sample one question from each of $k$ reasoning benchmarks to construct an $k$-problem prompt. In total, ZeMPE comprises 53,100 zero-shot multi-problem prompts containing classification and reasoning problems.
Classification-Related Tasks
@@ 3.3.3 ZeMPE Composition
For each classification benchmarks, we consider 5 task sizes and ensure that each task size has 100 distinct prompt instances: 5, 10, 20, 50, and 100 for single-text benchmarks and 3, 5, 10, 20, and 50 for text-pair benchmarks. To isolate the effect of task size, a larger task size only differs from a smaller one by having additional problems given a benchmark; and to isolate the effect of task, different MPE tasks share the same sets of problems given a task size and a benchmark. In total, this results in 13,500 prompts for classification-related MPE tasks.
Reasoning-Related Tasks
Besides vanilla zero-shot prompting, we also perform zero-shot-CoT following Kojima et al. (2022). [^3] In our early experiments, we found that zero-shot-CoT did not lead to different responses for the classification-related MPE tasks probably due to their novelty, so it was not used. Inspired by Cheng et al. (2023) as well as to control for the number of prompts generated, we consider smaller task sizes from 2 to 10. To ensure a reliable evaluation (e.g., sufficient parsable outputs), we increase the number of prompts from 100 to 300 for each benchmark/task size combination.
For each reasoning benchmark, we consider task sizes 2, 5, and 10 for MultiReasonSS. To robustly examine an LLM’s performance on mixed-source prompts, we create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks. For each benchmark combination, we consider the first 2, 4, 6, 8, and 10 benchmarks (also equals the respective task sizes) in the combination for MultiReasonMS. We also control the effects of task size and task with careful sampling, similar to what we did above. This results in 21,600 and 18,000 prompts for MultiReasonSS and MultiReasonMS, respectively.
4 Experiments
This section first describes the experimental setups and then reports and discusses the results.
4.1 LLMs and Evaluation Settings
@@
We evaluate 7 LLMs from 4 model families with greedy decoding for the four classification-related tasks: Vicuna (13B, Chiang et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), Llama-3 8B and 70B (Instruct, Meta, 2024), GPT-3.5, and GPT-4 (OpenAI, 2023). See Appendix A for the details about the LLMs used.
Given the consistent results we observed across LLMs with the classification-related tasks and for budget reasons, we only use two LLMs with greedy decoding, i.e., GPT-3.5 and Llama-3 70B. Since Llama-3 models tend to produce reasoning steps even when not instructed to do so, we only prompt GPT-3.5 with zero-shot-CoT.
Figure 1: Average accuracy of the 7 LLMs on the 4 classification-related task across task sizes for each benchmark.
Figure 2: Cost/Accuracy Ratio (lower is better) for the 7 LLMs on the 6 benchmarks for SingleClf (task size=1) and BatchClf (otherwise). We use the average (input + output) token count per classification as the proxy for the actual inference costs, calculated on the basis of the input and output tokens. The plot for MRPC is log-scale for the y-axis.
4.2 Performance Metric
@@
We measure the average per-problem accuracy (PPA) to unify the evaluation across all proposed tasks. PPA, defined in Equation 1, is the average accuracy of classifying $n$ problems in each prompt or, in the case of SelectOne, in each set of directly related prompts targeting different class labels.
@@
where $I(P_{i})$ is the inferred LLM-generated answer to the $i$th problem in the input prompt, $A_{i}$ is the ground truth, and $\delta(i,j)=1$ iff $i=j$ and 0 otherwise.
For SelectOne and SelectAll, $I(P_{i})$ is determined by considering the LLM’s assignments of indices to all class labels. Other than assigning an index to a wrong class label, there are two more error types. First, LLMs may assign an index with more than one class label, i.e., a contradiction error. Second, LLMs may assign no labels to an index at all, namely, a non-excluded middle error.
For MultiReasonMS prompts containing $k$ problems evenly sampled from $k$ benchmarks, we compute the expected PPA by averaging over the observed SPP performance for each benchmark.
@@
4.3 Classification-Related Results
In line with previous studies (Cheng et al., 2023; Lin et al., 2024) on few-shot MPP, we observe that while large language models (LLMs) demonstrate strong zero-shot classification capabilities and prompting with multiple problems can be cost-efficient, their performance degrades significantly when the same sets of problems are presented in a different format.
Although the BatchClf accuracy generally declines as the task size increases (Fig 1), all 7 LLMs achieve accuracy of at least 90% that of SingleClf across the benchmarks most of the time (see Table 5 in Appendix B). Overall, the SingleClf accuracy for the 7 LLMs on the 6 benchmarks is 75.5% and the BatchClf accuracy is 72.3%, a minor 3.2% absolute drop from the former. Interestingly, for SNLI almost all LLMs perform better in BatchClf than in SingleClf across all the task sizes (3-50) and GPT-4 consistently achieves a BatchClf accuracy near or better than the SingleClf accuracy under all conditions (see Fig 7 in Appendix B).
Zero-shot MPP can be cost-efficient.
Single-problem prompting can waste input tokens by redundantly repeating a shared task instruction. Multi-problem prompting reduces this redundancy, and this saving is larger the more problems are combined in a single prompt. Because performance tends to decline slowly as the number of tasks increases, this yields a favorable cost-accuracy ratio as the number of tasks increases (Fig 2). We only encountered two outliers, Vicuna on MRPC with task size 50 where the average input is 3,645 tokens and the context window is 4,096 tokens, and Mixtral 8x7B at $\geq$ 50 on AGNews. While it is of course up to downstream users to determine what cost-accuracy is right for them, this is likely beneficial for many use cases where similar prompts are repeated frequently.
To illustrate, we choose for each model/task combination the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair. We observe that MPP reduces substantial inference costs for all LLMs run on the 6 benchmarks, ranging from from 30.7% to 82.0% (see Fig 6 in Appendix B).